How many sets of three consecutive integers are there for which their sum does not equal three times the middle integer?
Answer by josmiceli(6786)

Looks like zero is the answer
consecutive integers = n, n+1, n+2
n + n + 1+ n+ 2 = 3(n + 1)
3n + 3 = 3n + 3
This is always true no matter what 3 consecutive integers you pick